Calculate mass of Al2(so4)3 in 250 ml of solution if the concentration is 0.48 N.
1
Expert's answer
2021-01-04T03:46:43-0500

Normality = no. of equivalent/volume of solution

0.48N = (No. of equivalent/0.25L)

no. of equivalent = 0.12

no. of moles / n-factor = 0.12

no. of moles = 0.12× 3 = 0.36

mass/molar mass = 0.36

mass = 0.36×342 = 123.12gram